OpenStudy (rainbow_dashie):

How many weeks will it take Justin to save $133 if he saves $9.50/wk? Complete the equation to model the problem. Use w to represent the unknown value. =

OpenStudy (rainbow_dashie):



OpenStudy (rainbow_dashie):

@iGreen @whatisthequestion @dan815 @TheSmartOne @Yolo4mecuite @UnkleRhaukus @Preetha @adrynicoleb @ganeshie8 @jose12345 @kl0723 @Lyrae @Callisto @Conqueror @NoelGreco

OpenStudy (anonymous):

$9.50 x W = $133 W = 133/9.50 W = 14 wk

OpenStudy (anonymous):

if u only need the equation then its this $9.50 x W = $133

OpenStudy (anonymous):

9.50 x week=133 Now you divide 133/9.50 133/9.50=14

OpenStudy (anonymous):

So 9.50*14=133

OpenStudy (anonymous):

@Rainbow_dashie

OpenStudy (rainbow_dashie):

ya

OpenStudy (rainbow_dashie):

is that right?

OpenStudy (rainbow_dashie):

only text allowed @CO_oLBoY

OpenStudy (rainbow_dashie):